A boy whose mass is 40kg rams up a flight of 30 steps each. 150mm high, in 6s. Find the average power developed
1
Expert's answer
2019-07-04T09:15:24-0400

We can find the average power developed from the formula:


"P = \dfrac{W}{t} = \dfrac{mgh}{t},""P = \dfrac{40 kg \cdot 10 \dfrac{m}{s^2} \cdot 30 steps \cdot 150 mm \cdot \dfrac{1m}{1000mm}}{6 s} = 300 W."

Answer:

"P = 300 W."